Recently, L. Chiatti Company paid a $2.37 dividend per share. The dividend is expected to grow at a constant rate of 5.50% forever. The common stock of the company currently sells for $52.50 per share. The before-tax cost of debt is 7.50%, and the company’s tax rate is 40%. The target capital structure of the company consists of 45% debt and 55% common equity. What is the company’s WACC if all the equity used is from retained earnings (Note, you must also explain your approach as to how to solve it in writing)?
Answer #1

Solution:

Dividend paid ( D0)= $2.37

Growth rate (g) = 5.50%

Next expected dividend ( D1) = $2.37(1+0.055)= $2.50

Price ( P0)= $52.50

Cost of equity = D1/P0 + g

Cost of equity = $2.50/$52.50 + 0.055= 10.26%

Before tax cost of debt =7.50%

Tax rate =40%

After tax cost of debt = 7.50%*(1-0.4)= 4.50%

WACC= cost of debt * weight of debt + cost of equity * weight of equity

WACC= 4.50%*0.45 + 10.26%*0.55

WACC= 2.025% + 5.643%= 7.668% rounded to 7.67%

WACC= 7.67%